### v1

# DISTA-Net: Dynamic Closely-Spaced Infrared Small Target Unmixing

Abstract
Resolving closely-spaced small targets in dense clusters presents a significant challenge in infrared imaging, as the overlapping signals hinder precise determination of their quantity, sub-pixel positions, and radiation intensities. While deep learning has advanced the field of infrared small target detection, its application to closely-spaced infrared small targets has not yet been explored. This gap exists primarily due to the complexity of separating superimposed characteristics and the lack of an open-source infrastructure. In this work, we propose the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net), which reconceptualizes traditional sparse reconstruction within a dynamic framework. DISTA-Net adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. To the best of our knowledge, DISTA-Net is the first deep learning model designed specifically for the unmixing of closely-spaced infrared small targets, achieving superior sub-pixel detection accuracy. Moreover, we have established the first open-source ecosystem to foster further research in this field. This ecosystem comprises three key components: (1) CSIST-100K, a publicly available benchmark dataset; (2) CSO-mAP, a custom evaluation metric for sub-pixel detection; and (3) GrokCSO, an open-source toolkit featuring DISTA-Net and other models. Our code and dataset are available at https://github.com/GrokCV/GrokCSO.
1 Introduction
Infrared imaging plays a pivotal role in various long-distance detection and surveillance tasks Sobrino et al. (2016), due to its exceptional sensitivity to thermal radiation and independence from illumination conditions. However, the radiation intensity captured from remote targets is inherently weak due to their long-range distance from the imaging system Kou et al. (2023). This challenge is exacerbated when targets appear in spatially close dense clusters, as the closely-spaced objects (CSO) Pruett et al. (2023), resulting in overlapping blob-like spots.
As shown in Fig. 1, such overlap makes it impossible to resolve the targets independently via human vision, thus obscuring the perception of target count, precise locations, and radiation intensities Zhao et al. (2022), presenting a substantial impediment for Infrared Search and Tracking (IRST) systems in their subsequent phases of detection, tracking, and identification. Therefore, exploring effective techniques for unmixing and reconstruction of such closely-spaced infrared small targets (CSIST), to accurately discern their exact locations and radiant intensities, holds great significance.
Figure 1: Conceptual illustration of imaging and unmixing processes for closely-spaced infrared small targets (CSIST). CSIST unmixing aims to disentangle and accurately estimate the count, positions, and intensities of overlapping targets.
Despite the critical importance of closely-spaced infrared small target unmixing in various applications, research addressing this specific task remains exceedingly scarce. These approaches typically formulate the problem as a parameter estimation task and employ optimization algorithms to solve it Meyer and Williams (2021). Target sparsity on the imaging plane was leveraged to devise a discretized sampling-based sparse reconstruction method Zhang et al. (2013), utilizing an over-complete dictionary and solving a second-order cone programming problem under the $\ell_{1}$ norm regularization.
However, the performance of these optimization-based models is highly dependent on meticulous hyperparameter tuning Chen et al. (2021), which poses significant challenges in real-world scenarios. Variations in target quantity or location further complicate the selection of optimal hyperparameters, limiting the generalizability and practicality of these methods Wan et al. (2018). Consequently, there is a pressing need for the development of unmixing algorithms that exhibit greater robustness to hyperparameter variations and can be effectively applied in diverse real-world settings.
While deep learning has revolutionized image super-resolution Li et al. (2023), its application to closely-spaced infrared small target unmixing remains unexplored, primarily due to fundamentally different task objectives and ecosystem limitations. Unlike generic super-resolution that enhances clarity through high-frequency detail restoration, infrared small target unmixing requires precise estimation of overlapping targets’ count, locations, and radiation intensities - essentially mapping images to specific target attributes rather than conventional pixel-space super-resolution. This challenge is further compounded by the absence of standardized benchmark datasets, task-specific evaluation metrics, and open-source implementations, creating significant barriers to developing and comparing deep learning approaches in this specialized domain.
To address the aforementioned challenges, in this paper, we propose a novel deep unfolding network for closely-spaced infrared small target unmixing, termed as the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net). We reformulate the traditional sparse reconstruction approach into a dynamic deep learning framework, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time. Distinct from prior methods, the parameters associated with the proximal mapping, including nonlinear transforms and shrinkage thresholds, are dynamically adapted to the input data, rather than being hand-crafted or fixed after training. To the best of our knowledge, this represents the first deep learning-based effort towards unmixing closely-spaced infrared small targets.
Furthermore, we have established a comprehensive open-source ecosystem to facilitate research in this domain. This ecosystem includes CSIST-100K (Closely-Spaced Infrared Small Target), an open benchmark dataset comprising 100,000 pairs of low-resolution images and exact annotations of location and radiation intensities. We also introduce CSO-mAP, a custom evaluation metric inspired by the mean average precision (mAP) from object detection, calibrated to evaluate the quantity, spatial positioning, and radiation intensity of the unmixed infrared small targets. Accompanying these resources is GrokCSO, an open-source PyTorch-based toolkit encapsulating our DISTA-Net alongside other state-of-the-art models, empowering researchers to effortlessly leverage the CSIST-100K dataset.
In summary, our contributions can be categorized into Four main aspects:
  1. 1.
    We reformulate the unmixing of closely-spaced infrared small targets as an interpretable deep unfolding problem. To our knowledge, this is the first deep learning based attempt for this task.
  2. 2.
    Our proposed DISTA-Net is a dynamic deep unfolding network, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process conditioned on the input data.
  3. 3.
    We establish the first open-source ecosystem for this task, including the CSIST-100K dataset, the CSO-mAP metric, and the GrokCSO toolkit.
  4. 4.
    We provide a comprehensive analysis of our approach, validating the importance of dynamic deep unfolding and the effectiveness of the DISTA-Net in addressing the CSIST unmixing task.
2 Relate Work
2.1 Infrared Small Target Detection
Driven by a range of open-source datasets Dai et al. (2021a, 2023b); Li et al. (2024b), infrared small target detection has garnered significant research attention in recent years. Current research mainly focuses on developing multi-scale feature fusion models to counteract the scarcity of intrinsic target features Xu et al. (2023). Dai et al. introduced an asymmetric contextual modulation module that bridges high-level semantics with low-level details using a bottom-up pathway via point-wise channel attention Dai et al. (2021b). Wang et al. merged reinforcement learning with pyramid feature fusion and proposed a global context boundary attention module to mitigate localized bright noise Wang et al. (2023). Tong et al. adopted an encoder–decoder structure, enhancing feature extraction with an atrous spatial pyramid pooling and a dual-attention module, and using multiscale labels to focus on target edges and internal features Tong et al. (2023).
Our work focuses on infrared small targets but differs from previous studies in two main ways. First, while infrared small target detection is a precursor to our study, our primary focus is on the unmixing of closely-spaced infrared small targets. Detecting potentially overlapped targets is essential for subsequent tasks like sub-pixel localization and radiation intensity prediction. Second, our task extends beyond simple detection by precisely locating targets at the sub-pixel level and estimating their radiative intensities, offering a more detailed understanding of target characteristics compared to the previous binary detection approach.
2.2 Deep Unfolding
Deep unfolding, as delineated in Monga et al. (2021), originated in 2010 with Learned Iterative Shrinkage-Thresholding Algorithm (LISTA) Gregor and LeCun (2010), which reinterprets ISTA Daubechies et al. (2004) as a fully connected feed-forward neural network. This approach generalizes effectively to new samples, achieving ISTA-like accuracy with fewer iterations. Subsequent works, such as ADMM-Net yang et al. (2016), have unfolded the steps of the Alternating Direction Method of Multipliers (ADMM) into a deep learning framework, thereby improving the accuracy and efficiency of MRI reconstruction through a compressive sensing model optimized via end-to-end discriminative training. Likewise, ISTA-Net Zhang and Ghanem (2018) has adopted an end-to-end learning approach for proximal mapping, enhancing the performance of compressive sensing for natural image reconstruction.
Motivated by such advances, deep unfolding has found applicability in a variety of computer vision tasks. Notably, Li et al. transformed a generalized gradient-domain total variation algorithm into a deep interpretable network for blind image deblurring, delivering superior performance with learned parameters Li et al. (2020). For image super-resolution, Guo et al. incorporated trainable convolutional layers into the Discrete Cosine Transform framework, effectively mitigating artifacts and enabling learning from limited data Guo et al. (2019). Solomon et al. unfolded robust principal component analysis into a deep network, improving the distinction between microbubble and tissue signals in ultrasound imaging Solomon et al. (2020).
Unlike previous methods with static parameters Zhang and Ghanem (2018); You et al. (2021), our DISTA-Net dynamically adapts proximal mapping weights based on the input, enabling an adaptive reconstruction process that caters to varying scenarios.
3 CSIST Benchmark, Metric, and Toolkit
Figure 2: CSIST Visualization: The top row shows 1 to 5 overlapping targets, and the following rows display unmixing results for sub-pixel division factors of 3×, 5×, and 7×.
CSIST-100K Dataset. When a single point source is imaged through a lens, an Airy pattern emerges at the focal point, characterized by a central bright spot comprising 84% of the total energy, surrounded by multiple concentric diffraction rings. This diffraction effect is typically approximated by the point spread function (PSF), whose standard deviation, dependent on the sensor’s focal ratio ($f$-number) and detection band, quantifies the energy spread. In multi-target imaging, the intensity captured by each pixel is the cumulative response of the overlaid point sources. Particularly for remote objects, each is considered a point source, with the radius of the resultant Airy spot dictated by $1.22\lambda/D$, where $\lambda$ is the wavelength and $D$ the lens diameter. This radius is equivalent to $1.9\sigma$ of a Gaussian PSF, marking the sensor’s physical resolution limit, as per the Rayleigh unit.
In our study, we set $\sigma_{\mathrm{PSF}}$ at 0.5 pixels. Our simulations include 1 to 5 targets per image, each characterized by sub-pixel coordinates and radiation intensity. Targets with radiation intensities between 220 and 250 units are randomly positioned within a single pixel in an $11\times 1$ 11 pixel grid, while maintaining separation of at least 0.52 Rayleigh units. We generated the CSIST-100K dataset with 100,000 samples, using 80,000 for training and dividing the remaining 20,000 equally between validation and testing. As shown in Fig. 2, closely-spaced targets undergo diffusion within the pixel matrix, with energy primarily concentrated in a $3\times 3$ pixel area. This creates significant overlap that makes it extremely difficult to determine both the number of distinct targets and their precise spatial coordinates.
CSO-mAP Metric.
Traditional bounding box metrics fails when targets lie within $1.9\sigma$ (due to Airy spot interference), Our CSO-mAP with strict sub-pixel position/intensity matching for CSIST evaluation, redefining TP/FP as follows:
$\mathbbm{1}_{k}\left(\hat{t}_{j},t_{i}\right)=\left\{\begin{array}[]{lc}1,&\text{ if }d\left(\hat{t}_{j},t_{i}\right)<\delta_{k}\\ 0,&\text{ otherwise }\end{array}\right.$ (1)
Here, $\hat{t}_{j}$ is the $j$-th predicted target, $t_{i}$ is the $i$-th true target, and $\delta_{k}\in\left\{0.05,0.1,0.15,0.2,0.25\right\}$ is a series of distance thresholds, with $k=1,2,3,4,5$, used to control the desired localization accuracy. Precision-recall (PR) curves are generated through intensity-prioritized matching, with Average Precision (AP) computed at each $\delta_{k}$. The final CSO-mAP metric is derived by averaging AP values across all thresholds, explicitly quantifying performance under varying spatial resolution demands, see Supplementary for details.
GrokCSO Toolkit. To address the lack of specialized tools in this domain, we introduce GrokCSO, an open-source toolkit for enhancing the reconstruction of closely-spaced infrared small targets. Built on PyTorch, GrokCSO provides pre-trained models, reproducibility scripts, and specialized evaluation metrics tailored for CSO challenges. The detailed architecture and features of this toolkit are provided in Supplementary.
4 Method
In this section, We introduce the imaging model for closely-spaced infrared small targets, traditional sparse reconstruction approaches, and the proposed DISTA-Net architecture.
4.1 Imaging and Unmixing Framework
CSIST Imaging Model.
Given the significant distance between targets and the infrared detector, targets can be approximated as point sources. The optical system’s diffraction spreads the energy across adjacent pixels, described by a two-dimensional Gaussian point spread function (PSF) Liaudat et al. (2021):
$p(x,y)=\frac{1}{2\pi\sigma_{\mathrm{PSF}}^{2}}\exp\left[-\frac{\left(x-x_{t}\right)^{2}+\left(y-y_{t}\right)^{2}}{2\sigma_{\mathrm{PSF}}^{2}}\right]\text{, }$ (2)
where $\sigma_{\mathrm{PSF}}$ represents the diffusion variance, and $(x_{i},y_{i})$ denote the target’s coordinates. For $N$ closely-spaced targets, each target is represented as $t_{i}=(x_{i},y_{i},g_{i},\sigma_{\mathrm{PSF}})$, where $i\in{1,..,N}$, with $(x_{i},y_{i})$ denoting sub-pixel coordinates and $g_{i}$ the radiation intensities. The collections of center points and intensities are denoted as $\mathcal{P}=\left\{p_{i}|i\in\{1,..,N\}\right\}$ and $\mathcal{G}=\left\{g_{i}|i\in\{1,..,N\}\right\}$. On an infrared focal plane of $U\times V$ pixels, each pixel integrates the PSF within its boundaries:
$g_{i,j}\left(x_{t},y_{t}\right)=\int_{x_{i,j}-1/2D}^{x_{i,j}+1/2D}\int_{y_{i,j}-1/2D}^{y_{i,j}+1/2D}p(x,y)\mathrm{d}x\mathrm{\leavevmode\nobreak\ d}y,$ (3)
where $(x_{i,j},y_{i,j})$ is the pixel’s center, $i=0,…,U,j=0,…,V-1$, with $D$ being the pixel width. The focal plane measurement model is vectorized as:
$\mathbf{z}$ $=\mathbf{G}(\mathbf{x},\mathbf{y})\mathbf{s}+\mathbf{n},$ (4)
where $\mathbf{G}(\mathbf{x},\mathbf{y})$ is the steering vector matrix, $\mathbf{s}$ represents target intensities, and $\mathbf{n}$ denotes Gaussian white noise.
CSIST Unmixing via Sparse Reconstruction. Given permissible quantization error, target positions in closely-spaced infrared small targets can be discretized into a finite set of sub-pixel locations $\Omega=\left\{(x_{l},y_{l})\right\}_{l=1,\cdots,L}$, where actual target positions form a sparse subset.
Figure 3: Division of each pixel into an $n\times n$ grid of sub-pixels, representing potential target positions.
As shown in Fig. 3, each pixel is divided into an $n\times n$ grid, resulting in $L=UVn^{2}$ sub-pixels. With sufficient grid resolution, each sub-pixel contains at most one target, with maximum position deviation of $\sqrt{2}D/n$. The measurement model can be extended to $\Omega$:
$\mathbf{z}=\mathbf{G}(\Omega)\tilde{\mathbf{s}}+\mathbf{w},$ (5)
where $\mathbf{G}(\Omega)$ contains steering vectors from $\Omega$, $\mathbf{w}$ denotes Gaussian white noise, and $\tilde{\mathbf{s}}\in\mathrm{R}^{L}$ is sparse with $L\gg UV$. The CSIST unmixing problem can then be formulated as a sparse reconstruction with $\ell_{1}$ regularization:
$\min_{\tilde{\mathbf{s}}}\|\mathbf{z}-\mathbf{G}(\Omega)\tilde{\mathbf{s}}\|_{2}^{2}+\lambda\|\tilde{\mathbf{s}}\|_{1},$ (6)
where $\lambda$ is the regularization parameter. The solution $\tilde{\mathbf{s}}$ directly yields target attributes: its non-zero entries indicate target count and intensities, while their corresponding positions in $\Omega$ provide sub-pixel coordinates.
Optimization Solution. The ISTA Daubechies et al. (2004) solves the sparse reconstruction problem through two alternating steps:
$\mathbf{r}^{(k)}$ $=\tilde{\mathbf{s}}^{(k-1)}-\rho\mathbf{G}^{\top}\left(\mathbf{G}\tilde{\mathbf{s}}^{(k-1)}-\mathbf{z}\right),$ (7)
$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\tilde{\mathbf{s}}-\mathbf{r}^{(k)}\right\|_{2}^{2}+\lambda\|\boldsymbol{\Psi}\tilde{\mathbf{s}}\|_{1}.$ (8)
where $\boldsymbol{\Psi}$ denotes the transform matrix, $k$ the iteration index, and $\rho$ the step size. The second step represents a proximal mapping:
$\operatorname{prox}_{\lambda\phi}(\mathbf{r})=\arg\min_{\tilde{\mathbf{s}}}\frac{1}{2}\|\tilde{\mathbf{s}}-\mathbf{r}\|_{2}^{2}+\lambda\phi(\tilde{\mathbf{s}}).$ (9)
While ISTA with orthogonal transforms (e.g., wavelets) has efficient solutions, it faces challenges with complex transforms and requires numerous iterations. To address these limitations, ISTA-Net replaces $\mathbf{\Psi}$ with a trainable non-linear transform $\mathcal{F}(\cdot)$:
$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\mathcal{F}\left(\tilde{\mathbf{s}}\right)-\mathcal{F}\left(\mathbf{r}^{(k)}\right)\right\|_{2}^{2}+\theta\|\mathcal{F}\left(\tilde{\mathbf{s}}\right)\|_{1},$ (10)
where $\theta$ is a learnable parameter. However, ISTA-Net’s static network weights post-training limit its adaptability to input data, particularly in CSIST unmixing scenarios where input sensitivity is crucial.
Figure 4: Architecture of the proposed DISTA-Net. The overall framework consists of multiple cascaded stages. Each stage contains three main components: a dual-branch dynamic transform module ($\mathcal{F}_{d}^{(k)}$) for feature extraction, a dynamic threshold module ($\Theta_{d}^{(k)}$) for feature refinement, and an inverse transform module ($\tilde{\mathcal{F}_{d}}^{(k)}$) for reconstruction.
4.2 DISTA-Net: A Dynamic Framework
To integrate ISTA with network-based sparse representation, DISTA-Net maps each ISTA iteration to a network layer. By stacking these layers and enabling end-to-end training, the network optimizes parameters automatically, improving upon traditional ISTA’s manually designed parameters while maintaining interpretability. Building upon ISTA-Net’s framework for compressed sensing, we introduce two key improvements. First, we design a data-adaptive nonlinear transformation function $\mathcal{F}_{d}(\cdot)$ that maps images to richer dimensional representations while emphasizing significant image regions. The transformation follows:
$\mathcal{F}_{d}(\tilde{\mathbf{s}}^{(k)})=\operatorname{Soft}(\mathcal{F}_{d}(\mathbf{r}^{(k)}),\theta_{d}),$ (11)
where $\operatorname{Soft}(\cdot,\theta_{d})$ denotes soft-thresholding with learnable parameter $\theta_{d}$, and $k$ is the stage index. The left inverse $\tilde{\mathcal{F}}$ satisfies ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with a structure symmetric to $\mathcal{F}_{d}(\cdot)$. This yields:
$\tilde{\mathbf{s}}^{(k)}=\tilde{\mathcal{F}}(\operatorname{Soft}(\mathcal{F}_{d}(\mathbf{r}^{(k)}),\theta_{d})).$ (12)
Second, we introduce a dynamic threshold module that adapts to input image variations, addressing the sensitivity of sparse vector perturbations in image generation. This flexible thresholding mechanism improves upon fixed parameters that can be either too strict or too lenient. As shown in Fig. 4, DISTA-Net comprises $N$ stages. Each stage $k$ ($k>1$) contains three components: $\mathcal{F}_{d}^{(k)}$, $\operatorname{Soft}(\cdot,\theta_{d}^{(k)})$, and ~$\tilde{\mathcal{F}}^{(k)}$. The input $\mathbf{r}^{(k)}$ is derived from $\tilde{\mathbf{s}}^{(k-1)}$ via Eq. (7), processed through these components sequentially to generate $\tilde{\mathbf{s}}^{(k)}$ for the next iteration.
Dynamic Transform.
While trainable non-linear transformations overcome limitations of handcrafted methods, their fixed post-training parameters result in static transformation patterns. We address this by introducing a dual-branch Dynamic Transform module $\mathcal{F}_{d}^{(k)}$ at the $k$-th stage (Fig. 4).
To handle the sensitivity of sparse image $\tilde{\mathbf{s}}^{(k-1)}$ perturbations, we design an auxiliary branch that guides $\mathbf{r}^{(k)}$ through a dynamic convolutional kernel. This approach enhances feature representation adaptively.
The module first processes $\tilde{\mathbf{s}}^{(k-1)}$ through a fully connected network to generate a weight vector:
$W=f(\tilde{\mathbf{s}}^{(k-1)})$ (13)
The Dynamic Conv module(Fig. 4) applies this weight vector as an adaptive convolutional kernel to $\mathbf{r}^{(k)}$:
$w_{r}=C(W,\mathbf{r}^{(k)})$ (14)
The final output combines a Conv-ReLU-Conv branch with the sigmoid-activated auxiliary branch:
$\mathcal{F}_{d}^{(k)}=\alpha\cdot A(\text{ReLU}(B(\mathbf{r}^{(k)})))+(1-\alpha)\cdot\text{sigmoid}(w_{r}),$ (15)
where $A(\cdot)$ and $B(\cdot)$ are convolution operations and $\alpha\in\left[0,1\right]$ governs the contribution between the two branches.
Dynamic Soft-Thresholding. Unlike ISTA-Net’s fixed threshold $\theta$, we propose a Dynamic Soft-Thresholding module that adapts $\theta_{d}$ based on input image information. This approach better handles densely overlapped targets and spatial context variations.
As shown in Fig. 4, the module employs dual convolutional layers to capture multi-scale features as in Li et al. (2024a, 2019a). The $\mathcal{F}_{d}^{(k)}$ output passes through two $3\times 3$ convolutions, generating feature maps $\tilde{U}_{1}$ and $\tilde{U}_{2}$. These are concatenated to form $\tilde{U}=[\tilde{U}_{1},\tilde{U}_{2}]$. Spatial relationships are captured through parallel pooling operations:
$SA_{\text{avg}}=P_{\text{avg}}(\tilde{U}),\quad SA_{\text{max}}=P_{\text{max}}(\tilde{U}).$ (16)
The pooled features are processed through the convolution:
$(\hat{SA})=\text{Conv}^{2\to N}([\text{SA}_{\text{avg}};\text{SA}_{\text{max}}]),$ (17)
followed by a sigmoid activation to generate spatial selective masks:
$(\tilde{SA})_{i}=\text{sigmoid}((\hat{SA})_{i}).$ (18)
The dynamic threshold $\theta_{d}$ is then computed by combining these masks with multi-scale feature maps through a final convolution.
Initialization and Learning Objectives.
DISTA-Net employs linear initialization similar to iterative sparse coding algorithms, where the initial estimate $\tilde{s}^{(0)}$ is obtained through an optimal linear projection $Q_{\text{init}}$ learned from training pairs $\{(z_{i},s_{i})\}$ (details in Supplementary).
The training objective combines reconstruction fidelity with structural preservation:
$\mathcal{L}=\mathcal{L}_{\text{discrepancy}}+\gamma\mathcal{L}_{\text{constraint}},$ (19)
where:
$\mathcal{L}_{\text{discrepancy}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\|\tilde{s}_{i}^{(N)}-s_{i}\|_{2}^{2},$ (20)
$\mathcal{L}_{\text{constraint}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\sum_{k=1}^{N}\|\tilde{\mathcal{F}}^{(k)}(\mathcal{F}_{d}^{(k)}(s_{i}))-s_{i}\|_{2}^{2}.$ (21)
Here, $\mathcal{L}_{\text{discrepancy}}$ computes the MSE between reconstructed $\tilde{s}_{i}^{(N)}$ and ground truth $s_{i}$, while $\mathcal{L}_{\text{constraint}}$ enforces multi-stage identity constraints through ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with $\gamma=0.01$ balancing these objectives.
5 Experiments
Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP PSNR $\uparrow$ SSIM $\uparrow$
mAP AP-05 AP-10 AP-15 AP-20 AP-25
                                          Traditional Optimization
ISTA Daubechies et al. (2004) - - 7.46 0.01 0.31 2.39 9.46 25.14 - -
Image Super-Resolution
ACTNet Zhang et al. (2023) 46.212M 62.80G 45.61 0.38 7.46 41.13 83.12 95.95 35.4526 99.70
CTNet Lim et al. (2017) 0.400M 2.756G 45.11 0.38 7.53 40.39 82.11 95.14 35.1499 99.70
DCTLSA Zeng et al. (2023) 0.865M 13.69G 44.51 0.39 7.35 39.35 81.15 94.34 34.6314 99.65
EDSR Lim et al. (2017) 1.552M 12.04G 45.32 0.33 7.07 40.58 83.24 95.41 35.3724 99.71
EGASR Qiu et al. (2023) 2.897M 17.73G 45.51 0.42 8.03 41.32 85.71 95.08 34.5681 99.66
FeNet Wang et al. (2022) 0.348M 2.578G 45.77 0.42 8.19 42.13 83.30 94.80 34.1531 99.66
RCAN Zhang et al. (2018a) 1.079M 8.243G 45.87 0.42 7.96 41.81 83.61 95.57 35.2119 99.69
RDN Zhang et al. (2018b) 22.306M 173.0G 45.81 0.35 7.11 41.07 84.07 96.43 36.4686 99.74
SAN Dai et al. (2019) 4.442M 34.05G 45.95 0.36 7.35 41.17 84.32 96.57 36.5037 99.74
SRCNN Dong et al. (2015) 0.019M 1.345G 29.06 0.23 4.10 21.65 49.95 69.39 28.7608 98.44
SRFBN Li et al. (2019b) 0.373M 3.217G 46.05 0.43 8.31 42.83 83.72 94.95 34.0174 99.68
HAN Niu et al. (2020) 64.342M 495.0G 45.70 0.39 7.46 40.90 83.61 96.17 35.2703 99.71
HiT-SIR Zhang et al. (2024) 0.769M 7.357G 44.51 0.40 7.16 39.55 80.94 94.50 34.7400 99.70
HiT-SNG Zhang et al. (2024) 0.952M 13.324G 45.01 0.39 7.34 40.19 81.98 95.17 35.1390 99.71
HiT-SRF Zhang et al. (2024) 0.844M 7.945G 44.71 0.40 7.71 40.03 80.94 94.49 34.5449 99.67
Deep Unfolding
ISTA-Net Zhang and Ghanem (2018) 0.171M 12.77G 45.16 0.41 7.71 40.57 82.58 94.53 33.9215 99.68
ISTA-Net+ Zhang and Ghanem (2018) 0.337M 24.33G 46.06 0.42 7.66 41.58 84.46 96.17 36.0892 99.72
LAMP Daubechies et al. (2004) 2.126M 0.278G 14.22 0.05 1.11 7.31 21.56 41.06 27.8299 96.89
LIHT Daubechies et al. (2004) 21.10M 1.358G 10.35 0.06 0.92 4.99 14.74 30.5 27.5107 96.42
LISTA Daubechies et al. (2004) 21.10M 1.358G 30.13 0.25 4.13 22.29 51.18 72.82 29.8936 99.12
FISTA-Net Xiang et al. (2021) 0.074M 18.96G 44.66 0.45 7.68 39.74 81.24 94.19 35.7519 99.67
TiLISTA Daubechies et al. (2004) 2.126M 0.278G 14.95 0.06 1.23 7.72 22.50 46.23 27.7038 97.40
$\star$ DISTA-Net (Ours) 2.179M 35.10G 46.74 0.38 7.54 42.44 86.18 97.14 37.8747 99.79
Table 1: Comparison with SOTA methods on the CSIST-100K dataset.
5.1 Experimental Settings
Training. Using images from the CSIST dataset as input, we employ the method mentioned in  4.1 to perform sub-pixel division with a sampling grid ratio of $c$, generating an unmixed high-resolution grid as ground truth. Specifically, for the $i$-th target with parameters $(x_{i},y_{i},g_{i})$, the intensity $g_{i}$ is assigned to the pixel at position $\left(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2}\right)$, with other pixels set to zero. With an appropriate selection of c, each target is resolved as a distinct point in the unmixed grid, enabling precise spatial separation.
Testing. We project the unmixed high-resolution grid back to the original 11×11 coordinate space to calculate the CSO-mAP against the ground truth target positions. Post-processing is applied to the model’s predicted images, where pixels with intensity values above a threshold of 50 are retained as predicted targets. For the i-th predicted target, its coordinates are computed as $\left(\frac{x_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c},\frac{y_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c}\right)$. To provide a more comprehensive assessment of the model’s performance, we also compute PSNR and SSIM metrics by directly comparing the predicted high-resolution images with the sub-pixel divided ground truth images.
Hyperparameters. We set $c=3$ as the baseline. The network is trained on randomly selected batches of 64 images, with the DIST-Net configured with $N=6$ stages and the Dynamic Transform branch coefficient $(1-\alpha)$ set to 0.3.
5.2 Comparison with SOTA Methods
Experimental Results. Table  1 compares DISTA-Net with traditional optimization (ISTA), image super-resolution, and deep unfolding methods on the CSIST-100K dataset, evaluated by computational efficiency (#P/FLOPs), localization accuracy (CSO-mAP), and image quality (PSNR/SSIM).
For localization accuracy, we adopt CSO-mAP with different distance thresholds (from AP-05 to AP-25), where a smaller threshold indicates stricter localization precision requirement. For instance, AP-05 evaluates the detection accuracy within 0.05-pixel distance, representing an extremely high precision demand. Under practical scenarios where 0.25-pixel precision (AP-25) is sufficient, DISTA-Net achieves 97.14% accuracy, outperforming most existing methods. Even under more stringent conditions (AP-15 and AP-10), our method maintains competitive performance with 42.44% and 7.54% accuracy respectively, demonstrating its robust localization capability across different precision requirements.
In terms of model efficiency, DISTA-Net achieves these results with moderate computational costs (2.179M parameters and 35.103G FLOPs), showing better efficiency compared to methods like ACTNet (46.212M, 62.798G) and HAN (64.342M, 0.495T). Additionally, DISTA-Net demonstrates superior image quality with the highest PSNR (37.8747) and SSIM (99.79) scores among all compared methods, indicating its excellent capability in preserving image details while performing target localization.
These comprehensive results validate that DISTA-Net achieves an effective balance between computational efficiency, localization accuracy, and reconstruction quality.
Figure 5: Visual comparison of 3× sub-pixel division reconstruction for scenes containing different numbers of adjacent targets. The red boxes highlight regions where targets exhibit significant sub-pixel characteristics.
Visual comparison. Fig. 5 compares reconstruction results with 3× sub-pixel division across methods for scenes containing 3–5 adjacent targets, assessing performance in dense multi-target scenarios. Existing methods struggle to reconstruct closely-spaced targets, exhibiting blurred boundaries or merged detections, with degradation worsening for higher target counts (e.g., five-target cases). In contrast, DISTA-Net preserves both target quantity and sub-pixel positions while maintaining sharp boundaries and accurate spatial distributions, even under extreme density.
5.3 Ablation Study
Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP
  ISTA-Net 0.171M 12.766G 45.16
DISTA-Net w/o DT 0.171M 39.544G 46.32
DISTA-Net w/o Thres. 5.111M 92.62G 46.17
$\star$ DISTA-Net (Ours) 2.179M 35.103G 46.74
Table 2: The effect of different components.
Effect of different components.
We conduct ablation studies to evaluate the contribution of each component in DISTA-Net, with results shown in Table 2. The second row (DISTA-Net w/o DT) corresponds to the model variant without Dynamic Transform and the third row (DISTA-Net w/o Thres.) represents the model without Dynamic Soft-Thresholding. Compared to the baseline ISTA-Net, our complete model shows notable improvements in both CSO-mAP (45.16% to 46.74%) and AP-20 (82.58% to 86.18%).
Removing the Dynamic Transform leads to a slight performance decrease (CSO-mAP drops to 46.32%), highlighting its role in enhancing the model’s performance. The removal of Dynamic Soft-Thresholding results in the most significant performance degradation (CSO-mAP decreases to 46.17%), emphasizing its key role in ensuring accuracy.
Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP
                                    c=5
ISTA-Net Zhang and Ghanem (2018) 0.171M 39.544G 66.90
ISTA-Net+ Zhang and Ghanem (2018) 0.225M 48.158G 68.50
CFGN Dai et al. (2023a) 0.538M 4.122G 67.95
$\star$ DISTA-Net (Ours) 5.153M 102.4G 69.58
c=7
ISTA-Net Zhang and Ghanem (2018) 0.171M 89.51G 71.19
ISTA-Net+ Zhang and Ghanem (2018) 0.225M 103.0G 71.09
CFGN Dai et al. (2023a) 0.548M 4.202G 70.38
$\star$ DISTA-Net (Ours) 6.409M 142.3G 72.84
Table 3: Comparison of methods across different Sampling Grids on the CSIST-100K dataset.
Model Performance and Sampling Grid. To investigate the impact of sampling grid ratio on model performance and computational efficiency, we conduct experiments with other different sampling grid ratios (c = 5, and 7) across various methods. Table 3 presents a comprehensive comparison of model performance under different sampling configurations.
As the sampling grid ratio increases, we observe a general trend of improved detection performance across all methods, which is attributed to the higher precision in target position representation. DISTA-Net consistently maintains a leading advantage in positioning accuracy, outperforming other methods at all configurations. However, this performance gain comes at the cost of substantially increased computational complexity. Notably, DISTA-Net demonstrates better efficiency in utilizing the increased resolution. Compared to ISTA-Net and ISTA-Net+, which show similar trends but lower performance, our method achieves higher accuracy gains with the same increase in sampling ratio. Based on these observations, we suggest selecting an appropriate sampling grid ratio according to real-world requirements to achieve a good balance between detection accuracy and computational efficiency.
Hyperparameters Analysis. We analyzed model performance versus stage numbers and dynamic branch coefficients in Supplementary, demonstrating robust architectural and parametric design.
6 Conclusion
In this work, we address the challenging problem of super-resolving closely-spaced infrared small targets, where accurate target counting, sub-pixel positioning, and radiation intensity measurement pose significant challenges. We present DISTA-Net, a deep learning framework that incorporates dynamic spatial attention mechanisms to reformulate traditional sparse reconstruction. Our comprehensive experimental results demonstrate that DISTA-Net achieves superior performance in both sub-pixel target detection accuracy and image reconstruction quality. To advance research in this domain, we introduce the CSIST dataset, CSO-mAP metric, and GrokCSO toolkit.
Supplementary
1 Multi-target Imaging
Figure 6: Multi target imaging, the imaging of distant targets on the image plane can be regarded as the diffusion of energy by point source targets through PSF function. Multi target imaging is the cumulative response of superimposed point sources.
When targets are at considerable distances, they can be approximated as point sources, with their energy distribution on the sensor’s focal plane characterized by Point Spread Functions (PSFs) that approximate two-dimensional Gaussian functions. When multiple targets are present in close proximity, the detector receives a combined energy response that is the linear sum of individual PSFs . This superposition principle is fundamental to the super-resolution unmixing process, where the stacked response must be decomposed to identify and characterize closely-spaced infrared targets, as illustrated in Fig. 6
2 CSO-mAP Metric
Once the predictions are classified as TP or FP, a binary list is constructed following the common practice in COCO, where a TP prediction is indicated by 1 and an FP prediction by 0. This binary list serves as the foundation for generating the Precision-Recall (PR) curve. By varying the confidence threshold for positive predictions, a series of Precision and Recall values are obtained, forming the PR curve. The Average Precision (AP), calculated as the area under the PR curve, provides a comprehensive assessment of the model’s performance across various confidence thresholds, effectively summarizing the trade-off between Precision and Recall. Finally, we introduce the CSO mean Average Precision (CSO-mAP), which averages the AP across different distance thresholds $\delta_{k}$, offering a standardized metric for comparing the performance of models in the CSIST unmixing task.
3 GrokCSO Toolkit
Recognizing the void in specialized, accessible tools for this domain, we introduce GrokCSO, a comprehensive, open-source toolkit architected to enhance the reconstruction of closely spaced infrared small targets. While generic computer vision fields benefit from a plethora of object detection toolkits like MMDetection and GluonCV, the specificity of infrared small target unmixing has long necessitated a dedicated platform. The absence of such a platform has fragmented research efforts, impeding the reproducibility of experiments and the comparative analysis of algorithms.
Developed atop the robust PyTorch framework, GrokCSO is meticulously designed to address the unique challenges inherent to closely spaced infrared small target unmixing. It stands distinguished in its offerings:
  1. 1.
    Pre-trained Models and Reproducibility: GrokCSO equips researchers with an arsenal of pre-trained models, complete with training scripts and logs of leading-edge algorithms. This repository of resources not only fosters reproducibility but also facilitates nuanced comparisons across diverse algorithmic approaches.
  2. 2.
    Tailored Flexibility and Evaluative Rigor: With an extensive suite of adaptable backbones and necks, GrokCSO accommodates a broader spectrum of computational strategies. The toolkit incorporates specialized dataset loaders, cutting-edge attention mechanisms, and versatile data augmentation workflows. Moreover, GrokCSO introduces evaluation metrics specifically calibrated for the intricacies of closely spaced infrared small target unmixing. These metrics honor the distinct nature of CSO challenges.
4 Hyperparameters Analysis
Figure 7: CSO-mAP performance under different Stage Numbers and Branch coefficients.
We also conduct experiments to analyze the relationship between model performance, the number of stages, and the coefficient of the dynamic transform branch in Supplementary, validating its architectural and parametric robustness.
First, to determine an appropriate number of stages, we trained the network with stage numbers $K=2,4,6,8,$ and $10$. The performance curves are recorded in Figure 7. As the number of stages increases from 2 to 6, the CSO-mAP score improves, reaching its peak (46.74%) at 6 stages. Beyond this point, additional stages result in diminishing returns, likely due to increased model complexity. Therefore, selecting 6 stages provides an optimal balance between accuracy and efficiency.
Second, we adjusted the proportion of the dynamic transform branch relative to the total contribution of both branches and observed the performance changes. Starting from 0% coefficient, performance improves as the influence of the dynamic branch increases, peaking at a 30% coefficient. The model maintains stable performance across coefficient settings from 50% to 90%, demonstrating robustness. However, at 100% coefficient, performance drops sharply to 30.62 (not labeled in the figure), confirming that integrating both branches is crucial for optimal feature representation.
5 Initialization and Learning Objectives.
Initialization. Given a dataset $\{(z_{i},s_{i})\}_{i=1}^{M}$, where $z_{i}$ is an image of overlapping targets and $s_{i}$ is the corresponding super-resolved image, we calculate high-resolution target coordinates in $s_{i}$ using the target information $(x_{i},y_{i},g_{i})$ and a scaling factor $c$. For instance, the grayscale value $g_{i}$ at position $(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2})$ in the high-resolution image is used to generate $s_{i}$.
Let $Z=[z_{1},\ldots,z_{M}]$ and $S=[s_{1},\ldots,s_{M}]$. The matrix $Q_{\text{init}}$ for initializing $\tilde{s}^{(0)}$ is computed as:
$Q_{\text{init}}=\arg\min_{Q}\|QZ-S\|_{F}^{2}=SZ^{T}(ZZ^{T})^{-1}.$
Thus, $\tilde{s}^{(0)}$ is given by:
$\tilde{s}^{(0)}=Q_{\text{init}}Z.$
Loss Function. To ensure that the unmixed image $\tilde{s}$ is as close as possible to the ground truth image $s$ while maintaining the structure $\tilde{F}(\cdot)\circ F(\cdot)=\mathbf{I}$, we designed the following end-to-end training loss function for DISTA-Net, with a training dataset of size $M$, $N$ stages, and image size $N_{s}$:
$\mathcal{L}=\mathcal{L}_{\text{discrepancy}}+\gamma\mathcal{L}_{\text{constraint}},$
where:
$\mathcal{L}_{\text{discrepancy}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\|\tilde{s}_{i}^{(N)}-s_{i}\|_{2}^{2},$
$\mathcal{L}_{\text{constraint}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\sum_{k=1}^{N}\|\tilde{F}^{(k)}(F^{(k)}(s_{i}))-s_{i}\|_{2}^{2}.$
Here, $L_{\text{discrepancy}}$ measures the mean squared error (MSE) between the super-resolved image $\tilde{s}_{i}^{(N)}$ and the ground truth image $s_{i}$. $L_{\text{constraint}}$ enforces the structural constraint by ensuring that the composition of $\tilde{F}^{(k)}$ and $F^{(k)}$ approximates the identity transformation for each stage $k$. $\gamma$ is a parameter that balances the discrepancy and constraint terms.
This loss function aims to balance the accuracy of the unmixed images with the structural integrity of the transformation functions.

### v2

@@
 Moreover, we have established the first open-source ecosystem to foster further research in this field.
-This ecosystem comprises three key components: (1) CSIST-100K, a publicly available benchmark dataset; (2) CSO-mAP, a custom evaluation metric for sub-pixel detection; and (3) GrokCSO, an open-source toolkit featuring DISTA-Net and other models.
-Our code and dataset are available at https://github.com/GrokCV/GrokCSO.
+This ecosystem comprises three key components: (1) CSIST-100K, a publicly available benchmark dataset; (2) CSO-mAP, a custom evaluation metric for sub-pixel detection; and (3) GrokCSO, an open-source toolkit featuring DISTA-Net and other state-of-the-art models, available at https://github.com/GrokCV/GrokCSO.
 1 Introduction
-Infrared imaging plays a pivotal role in various long-distance detection and surveillance tasks Sobrino et al.
-(2016), due to its exceptional sensitivity to thermal radiation and independence from illumination conditions.
-However, the radiation intensity captured from remote targets is inherently weak due to their long-range distance from the imaging system Kou et al.
-(2023).
-This challenge is exacerbated when targets appear in spatially close dense clusters, as the closely-spaced objects (CSO) Pruett et al.
-(2023), resulting in overlapping blob-like spots.
+Infrared imaging plays a pivotal role in various long-distance detection and surveillance tasks [26], due to its exceptional sensitivity to thermal radiation and independence from illumination conditions.
+However, the radiation intensity captured from remote targets is inherently weak due to their long-range distance from the imaging system [11].
+This challenge is exacerbated when targets appear in spatially close dense clusters, as the closely-spaced objects (CSO) [24], resulting in overlapping blob-like spots.
 As shown in Fig.
-1, such overlap makes it impossible to resolve the targets independently via human vision, thus obscuring the perception of target count, precise locations, and radiation intensities Zhao et al.
-(2022), presenting a substantial impediment for Infrared Search and Tracking (IRST) systems in their subsequent phases of detection, tracking, and identification.
+1, such overlap makes it impossible to resolve the targets independently via human vision, thus obscuring the perception of target count, precise locations, and radiation intensities [46], presenting a substantial impediment for Infrared Search and Tracking (IRST) systems in their subsequent phases of detection, tracking, and identification.
 Therefore, exploring effective techniques for unmixing and reconstruction of such closely-spaced infrared small targets (CSIST), to accurately discern their exact locations and radiant intensities, holds great significance.
@@
 CSIST unmixing aims to disentangle and accurately estimate the count, positions, and intensities of overlapping targets.
-Despite the critical importance of closely-spaced infrared small target unmixing in various applications, research addressing this specific task remains exceedingly scarce.
-These approaches typically formulate the problem as a parameter estimation task and employ optimization algorithms to solve it Meyer and Williams (2021).
-Target sparsity on the imaging plane was leveraged to devise a discretized sampling-based sparse reconstruction method Zhang et al.
-(2013), utilizing an over-complete dictionary and solving a second-order cone programming problem under the $\ell_{1}$ norm regularization.
-However, the performance of these optimization-based models is highly dependent on meticulous hyperparameter tuning Chen et al.
-(2021), which poses significant challenges in real-world scenarios.
-Variations in target quantity or location further complicate the selection of optimal hyperparameters, limiting the generalizability and practicality of these methods Wan et al.
-(2018).
+Despite the critical importance of CSIST unmixing in various applications, research addressing this specific task remains exceedingly scarce.
+These approaches typically formulate the problem as a parameter estimation task and employ optimization algorithms to solve it [21].
+Target sparsity on the imaging plane was leveraged to devise a discretized sampling-based sparse reconstruction method [39], utilizing an over-complete dictionary and solving a second-order cone programming problem under the $\ell_{1}$ norm regularization.
+However, the performance of these optimization-based models is highly dependent on meticulous hyperparameter tuning [1], which poses significant challenges in real-world scenarios.
+Variations in target quantity or location further complicate the selection of optimal hyperparameters, limiting the generalizability and practicality of these methods [29].
 Consequently, there is a pressing need for the development of unmixing algorithms that exhibit greater robustness to hyperparameter variations and can be effectively applied in diverse real-world settings.
-While deep learning has revolutionized image super-resolution Li et al.
-(2023), its application to closely-spaced infrared small target unmixing remains unexplored, primarily due to fundamentally different task objectives and ecosystem limitations.
-Unlike generic super-resolution that enhances clarity through high-frequency detail restoration, infrared small target unmixing requires precise estimation of overlapping targets’ count, locations, and radiation intensities - essentially mapping images to specific target attributes rather than conventional pixel-space super-resolution.
+While deep learning has revolutionized image super-resolution [13], its application to CSIST unmixing remains unexplored, primarily due to fundamentally different task objectives and ecosystem limitations.
+Unlike generic super-resolution that enhances clarity through high-frequency detail restoration [33], infrared small target unmixing requires precise estimation of overlapping targets’ counts, locations, and radiation intensities - essentially mapping images to specific target attributes rather than conventional pixel-space super-resolution.
 This challenge is further compounded by the absence of standardized benchmark datasets, task-specific evaluation metrics, and open-source implementations, creating significant barriers to developing and comparing deep learning approaches in this specialized domain.
-To address the aforementioned challenges, in this paper, we propose a novel deep unfolding network for closely-spaced infrared small target unmixing, termed as the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net).
+To address the aforementioned challenges, in this paper, we propose a novel deep unfolding network for CSIST unmixing, termed as the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net).
 We reformulate the traditional sparse reconstruction approach into a dynamic deep learning framework, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process in real time.
-Distinct from prior methods, the parameters associated with the proximal mapping, including nonlinear transforms and shrinkage thresholds, are dynamically adapted to the input data, rather than being hand-crafted or fixed after training.
-To the best of our knowledge, this represents the first deep learning-based effort towards unmixing closely-spaced infrared small targets.
-Furthermore, we have established a comprehensive open-source ecosystem to facilitate research in this domain.
-This ecosystem includes CSIST-100K (Closely-Spaced Infrared Small Target), an open benchmark dataset comprising 100,000 pairs of low-resolution images and exact annotations of location and radiation intensities.
-We also introduce CSO-mAP, a custom evaluation metric inspired by the mean average precision (mAP) from object detection, calibrated to evaluate the quantity, spatial positioning, and radiation intensity of the unmixed infrared small targets.
-Accompanying these resources is GrokCSO, an open-source PyTorch-based toolkit encapsulating our DISTA-Net alongside other state-of-the-art models, empowering researchers to effortlessly leverage the CSIST-100K dataset.
-In summary, our contributions can be categorized into Four main aspects:
-  1. 1.
-    We reformulate the unmixing of closely-spaced infrared small targets as an interpretable deep unfolding problem. To our knowledge, this is the first deep learning based attempt for this task.
-  2. 2.
+Distinct from prior methods, the parameters associated with the proximal mapping (nonlinear transforms and shrinkage thresholds) are dynamically adapted to the input data, rather than being hand-crafted or fixed after training.
+To the best of our knowledge, this represents the first deep learning-based effort towards CSIST unmixing.
+Furthermore, we establish a comprehensive open-source ecosystem to facilitate research in this domain, including CSIST-100K, an open benchmark dataset comprising 100,000 pairs of CSIST images and exact annotations of location and radiation intensities; CSO-mAP, a custom evaluation metric inspired by the mean average precision (mAP) from object detection, calibrated to evaluate the quantity, spatial positioning, and radiation intensity of the unmixed infrared targets; and GrokCSO, an open-source PyTorch-based toolkit encapsulating our DISTA-Net alongside other state-of-the-art models, empowering researchers to effortlessly leverage the CSIST-100K dataset.
+Our contributions can be categorized into Four aspects: 1.
+We reformulate CSIST unmixing as an interpretable deep unfolding problem.
+To our knowledge, this is the first deep learning based attempt for this task.
+2.
 Our proposed DISTA-Net is a dynamic deep unfolding network, which adaptively generates convolution weights and thresholding parameters to tailor the reconstruction process conditioned on the input data.
-  3. 3.
+3.
 We establish the first open-source ecosystem for this task, including the CSIST-100K dataset, the CSO-mAP metric, and the GrokCSO toolkit.
-  4. 4.
+4.
 We provide a comprehensive analysis of our approach, validating the importance of dynamic deep unfolding and the effectiveness of the DISTA-Net in addressing the CSIST unmixing task.
-2 Relate Work
+2 Related Work
 2.1 Infrared Small Target Detection
-Driven by a range of open-source datasets Dai et al.
-(2021a, 2023b); Li et al.
-(2024b), infrared small target detection has garnered significant research attention in recent years.
-Current research mainly focuses on developing multi-scale feature fusion models to counteract the scarcity of intrinsic target features Xu et al.
-(2023).
+Driven by a range of open-source datasets [4, 6, 15], infrared small target detection has garnered significant research attention in recent years.
+Current research mainly focuses on developing multi-scale feature fusion models to counteract the scarcity of intrinsic target features [35].
 Dai et al.
-introduced an asymmetric contextual modulation module that bridges high-level semantics with low-level details using a bottom-up pathway via point-wise channel attention Dai et al.
-(2021b).
+introduced an asymmetric contextual modulation module that bridges high-level semantics with low-level details using a bottom-up pathway via point-wise channel attention [5].
 Wang et al.
-merged reinforcement learning with pyramid feature fusion and proposed a global context boundary attention module to mitigate localized bright noise Wang et al.
-(2023).
+merged reinforcement learning with pyramid feature fusion and proposed a global context boundary attention module to mitigate localized bright noise [32].
+Cheng et al.
+proposed a difference-aware attention module with a dual-temporal aggregation module for global feature learning and channel activation, and a difference-attention module for multi-scale detection via local correlations [20].
 Tong et al.
-adopted an encoder–decoder structure, enhancing feature extraction with an atrous spatial pyramid pooling and a dual-attention module, and using multiscale labels to focus on target edges and internal features Tong et al.
-(2023).
-Our work focuses on infrared small targets but differs from previous studies in two main ways.
-First, while infrared small target detection is a precursor to our study, our primary focus is on the unmixing of closely-spaced infrared small targets.
-Detecting potentially overlapped targets is essential for subsequent tasks like sub-pixel localization and radiation intensity prediction.
-Second, our task extends beyond simple detection by precisely locating targets at the sub-pixel level and estimating their radiative intensities, offering a more detailed understanding of target characteristics compared to the previous binary detection approach.
+adopted an encoder–decoder structure, enhancing feature extraction with an atrous spatial pyramid pooling and a dual-attention module, and using multiscale labels to focus on target edges and internal features [28].
+Zhang et al.
+proposed an infrared small target detection framework integrating visual-textual information via CLIP-prompted SAM adaptation with a denoising module, achieving enhanced generalization capability for the infrared domain [41].
+Our work focuses on infrared small targets but differs in two key aspects.
+First, while infrared small target detection precedes our study, we prioritize CSIST unmixing, where detecting overlapped targets is crucial for sub-pixel localization and radiation intensity prediction.
+Second, our task goes beyond detection by enabling sub-pixel-level localization and radiative intensity estimation, providing finer target characterization than binary detection.
 2.2 Deep Unfolding
-Deep unfolding, as delineated in Monga et al.
-(2021), originated in 2010 with Learned Iterative Shrinkage-Thresholding Algorithm (LISTA) Gregor and LeCun (2010), which reinterprets ISTA Daubechies et al.
-(2004) as a fully connected feed-forward neural network.
+Deep unfolding, as delineated in [22], originated in 2010 with the Learned Iterative Shrinkage-Thresholding Algorithm (LISTA) [9], which reinterprets ISTA [7] as a fully connected feed-forward neural network.
 This approach generalizes effectively to new samples, achieving ISTA-like accuracy with fewer iterations.
-Subsequent works, such as ADMM-Net yang et al.
-(2016), have unfolded the steps of the Alternating Direction Method of Multipliers (ADMM) into a deep learning framework, thereby improving the accuracy and efficiency of MRI reconstruction through a compressive sensing model optimized via end-to-end discriminative training.
-Likewise, ISTA-Net Zhang and Ghanem (2018) has adopted an end-to-end learning approach for proximal mapping, enhancing the performance of compressive sensing for natural image reconstruction.
+Subsequent works, such as ADMM-Net [36], have unfolded the steps of the Alternating Direction Method of Multipliers (ADMM) into a deep learning framework, thereby improving the accuracy and efficiency of MRI reconstruction through a compressive sensing model optimized via end-to-end discriminative training.
+Likewise, ISTA-Net [40] has adopted an end-to-end learning approach for proximal mapping, enhancing the performance of compressive sensing for natural image reconstruction.
 Motivated by such advances, deep unfolding has found applicability in a variety of computer vision tasks.
 Notably, Li et al.
-transformed a generalized gradient-domain total variation algorithm into a deep interpretable network for blind image deblurring, delivering superior performance with learned parameters Li et al.
-(2020).
+transformed a generalized gradient-domain total variation algorithm into a deep interpretable network for blind image deblurring, delivering superior performance with learned parameters [14].
 For image super-resolution, Guo et al.
-incorporated trainable convolutional layers into the Discrete Cosine Transform framework, effectively mitigating artifacts and enabling learning from limited data Guo et al.
-(2019).
+incorporated trainable convolutional layers into the Discrete Cosine Transform framework, effectively mitigating artifacts and enabling learning from limited data [10].
 Solomon et al.
-unfolded robust principal component analysis into a deep network, improving the distinction between microbubble and tissue signals in ultrasound imaging Solomon et al.
-(2020).
-Unlike previous methods with static parameters Zhang and Ghanem (2018); You et al.
-(2021), our DISTA-Net dynamically adapts proximal mapping weights based on the input, enabling an adaptive reconstruction process that caters to varying scenarios.
+unfolded robust principal component analysis into a deep network, improving the distinction between microbubble and tissue signals in ultrasound imaging [27].
+Unlike previous methods with static parameters [40, 37], our DISTA-Net dynamically adapts proximal mapping weights based on the input, enabling an adaptive reconstruction process that caters to varying scenarios.
 3 CSIST Benchmark, Metric, and Toolkit
-Figure 2: CSIST Visualization: The top row shows 1 to 5 overlapping targets, and the following rows display unmixing results for sub-pixel division factors of 3×, 5×, and 7×.
+Figure 2: CSIST Visualization: The top row shows 1 to 5 overlapping targets, and the following rows display unmixing results for sub-pixel division factors of $3\times$, $5\times$, and $7\times$.
 CSIST-100K Dataset.
-When a single point source is imaged through a lens, an Airy pattern emerges at the focal point, characterized by a central bright spot comprising 84% of the total energy, surrounded by multiple concentric diffraction rings.
-This diffraction effect is typically approximated by the point spread function (PSF), whose standard deviation, dependent on the sensor’s focal ratio ($f$-number) and detection band, quantifies the energy spread.
-In multi-target imaging, the intensity captured by each pixel is the cumulative response of the overlaid point sources.
-Particularly for remote objects, each is considered a point source, with the radius of the resultant Airy spot dictated by $1.22\lambda/D$, where $\lambda$ is the wavelength and $D$ the lens diameter.
-This radius is equivalent to $1.9\sigma$ of a Gaussian PSF, marking the sensor’s physical resolution limit, as per the Rayleigh unit.
 In our study, we set $\sigma_{\mathrm{PSF}}$ at 0.5 pixels.
-Our simulations include 1 to 5 targets per image, each characterized by sub-pixel coordinates and radiation intensity.
-Targets with radiation intensities between 220 and 250 units are randomly positioned within a single pixel in an $11\times 1$ 11 pixel grid, while maintaining separation of at least 0.52 Rayleigh units.
-We generated the CSIST-100K dataset with 100,000 samples, using 80,000 for training and dividing the remaining 20,000 equally between validation and testing.
+Simulations include 1–5 overlapping targets per image, each defined by 2D coordinates and radiation intensity (220–250 units), randomly placed within an $11\times 11$ grid while maintaining $\geq$ 0.52 Rayleigh units separation.
+We generated the CSIST-100K dataset with 100,000 samples: 80,000 for training, and 20,000 split equally for validation and testing.
 As shown in Fig.
-2, closely-spaced targets undergo diffusion within the pixel matrix, with energy primarily concentrated in a $3\times 3$ pixel area.
-This creates significant overlap that makes it extremely difficult to determine both the number of distinct targets and their precise spatial coordinates.
+2, closely spaced targets diffuse, with energy concentrated in a $3\times 3$ pixel area, causing significant overlap that complicates target counting and coordinate determination.
+(See Supplementary for optical modeling details.)
 CSO-mAP Metric.
-Traditional bounding box metrics fails when targets lie within $1.9\sigma$ (due to Airy spot interference), Our CSO-mAP with strict sub-pixel position/intensity matching for CSIST evaluation, redefining TP/FP as follows:
-$\mathbbm{1}_{k}\left(\hat{t}_{j},t_{i}\right)=\left\{\begin{array}[]{lc}1,&\text{ if }d\left(\hat{t}_{j},t_{i}\right)<\delta_{k}\\ 0,&\text{ otherwise }\end{array}\right.$ (1)
-Here, $\hat{t}_{j}$ is the $j$-th predicted target, $t_{i}$ is the $i$-th true target, and $\delta_{k}\in\left\{0.05,0.1,0.15,0.2,0.25\right\}$ is a series of distance thresholds, with $k=1,2,3,4,5$, used to control the desired localization accuracy.
+Traditional bounding box metrics fail when target separation falls below the Rayleigh criterion (due to Airy spot interference).
+Our CSO-mAP with strict sub-pixel position/intensity matching for CSIST evaluation, redefining TP/FP as follows:
+[…]
+Here, $\delta_{k}\in\left\{0.05,0.1,0.15,0.2,0.25\right\}$ is a series of distance thresholds, with $k=1,2,3,4,5$, used to control the desired localization accuracy.
 Precision-recall (PR) curves are generated through intensity-prioritized matching, with Average Precision (AP) computed at each $\delta_{k}$.
-The final CSO-mAP metric is derived by averaging AP values across all thresholds, explicitly quantifying performance under varying spatial resolution demands, see Supplementary for details.
+The final CSO-mAP metric is derived by averaging AP values across all thresholds, explicitly quantifying performance under varying spatial resolution demands (details in Supplementary).
 GrokCSO Toolkit.
-To address the lack of specialized tools in this domain, we introduce GrokCSO, an open-source toolkit for enhancing the reconstruction of closely-spaced infrared small targets.
+To address the lack of specialized tools in this domain, we introduce GrokCSO, an open-source toolkit for CSIST unmixing.
 Built on PyTorch, GrokCSO provides pre-trained models, reproducibility scripts, and specialized evaluation metrics tailored for CSO challenges.
@@
 Given the significant distance between targets and the infrared detector, targets can be approximated as point sources.
-The optical system’s diffraction spreads the energy across adjacent pixels, described by a two-dimensional Gaussian point spread function (PSF) Liaudat et al.
-(2021):
+The optical system’s diffraction spreads the energy across adjacent pixels, described by a two-dimensional Gaussian point spread function (PSF) [18]:
 $p(x,y)=\frac{1}{2\pi\sigma_{\mathrm{PSF}}^{2}}\exp\left[-\frac{\left(x-x_{t}\right)^{2}+\left(y-y_{t}\right)^{2}}{2\sigma_{\mathrm{PSF}}^{2}}\right]\text{, }$ (2)
-where $\sigma_{\mathrm{PSF}}$ represents the diffusion variance, and $(x_{i},y_{i})$ denote the target’s coordinates.
-For $N$ closely-spaced targets, each target is represented as $t_{i}=(x_{i},y_{i},g_{i},\sigma_{\mathrm{PSF}})$, where $i\in{1,..,N}$, with $(x_{i},y_{i})$ denoting sub-pixel coordinates and $g_{i}$ the radiation intensities.
-The collections of center points and intensities are denoted as $\mathcal{P}=\left\{p_{i}|i\in\{1,..,N\}\right\}$ and $\mathcal{G}=\left\{g_{i}|i\in\{1,..,N\}\right\}$.
+where $\sigma_{\mathrm{PSF}}^{2}$ is the diffusion variance and $(x_{t},y_{t})$ the target coordinates.
 On an infrared focal plane of $U\times V$ pixels, each pixel integrates the PSF within its boundaries:
-$g_{i,j}\left(x_{t},y_{t}\right)=\int_{x_{i,j}-1/2D}^{x_{i,j}+1/2D}\int_{y_{i,j}-1/2D}^{y_{i,j}+1/2D}p(x,y)\mathrm{d}x\mathrm{\leavevmode\nobreak\ d}y,$ (3)
-where $(x_{i,j},y_{i,j})$ is the pixel’s center, $i=0,…,U,j=0,…,V-1$, with $D$ being the pixel width. The focal plane measurement model is vectorized as:
+$g_{i,j}\left(x_{t},y_{t}\right)=\int_{x_{i,j}-1/2D}^{x_{i,j}+1/2D}\int_{y_{i,j}-1/2D}^{y_{i,j}+1/2D}p(x,y)\mathrm{d}x\mathrm{~d}y,$ (3)
+where $(x_{i,j},y_{i,j})$ is the pixel’s center and $D$ the pixel width. The focal plane measurement model is vectorized as:
 $\mathbf{z}$ $=\mathbf{G}(\mathbf{x},\mathbf{y})\mathbf{s}+\mathbf{n},$ (4)
-where $\mathbf{G}(\mathbf{x},\mathbf{y})$ is the steering vector matrix, $\mathbf{s}$ represents target intensities, and $\mathbf{n}$ denotes Gaussian white noise.
+where $\mathbf{G}(\mathbf{x},\mathbf{y})$ is the steering matrix, $\mathbf{s}$ represents target intensities, and $\mathbf{n}$ denotes Gaussian white noise.
 CSIST Unmixing via Sparse Reconstruction.
@@
 3, each pixel is divided into an $n\times n$ grid, resulting in $L=UVn^{2}$ sub-pixels.
-With sufficient grid resolution, each sub-pixel contains at most one target, with maximum position deviation of $\sqrt{2}D/n$.
+With sufficient grid resolution, each sub-pixel contains at most one target, with maximum position deviation of $\sqrt{2}D/2n$.
 The measurement model can be extended to $\Omega$:
 $\mathbf{z}=\mathbf{G}(\Omega)\tilde{\mathbf{s}}+\mathbf{w},$ (5)
-where $\mathbf{G}(\Omega)$ contains steering vectors from $\Omega$, $\mathbf{w}$ denotes Gaussian white noise, and $\tilde{\mathbf{s}}\in\mathrm{R}^{L}$ is sparse with $L\gg UV$.
+where $\mathbf{G}(\Omega)$ contains steering vectors from $\Omega$, and $\tilde{\mathbf{s}}\in\mathrm{R}^{L}$ is sparse with $L\gg UV$.
 The CSIST unmixing problem can then be formulated as a sparse reconstruction with $\ell_{1}$ regularization:
@@
 The solution $\tilde{\mathbf{s}}$ directly yields target attributes: its non-zero entries indicate target count and intensities, while their corresponding positions in $\Omega$ provide sub-pixel coordinates.
-Optimization Solution. The ISTA Daubechies et al. (2004) solves the sparse reconstruction problem through two alternating steps:
+Optimization Solution. The ISTA [7] solves the sparse reconstruction problem through two alternating steps:
 $\mathbf{r}^{(k)}$ $=\tilde{\mathbf{s}}^{(k-1)}-\rho\mathbf{G}^{\top}\left(\mathbf{G}\tilde{\mathbf{s}}^{(k-1)}-\mathbf{z}\right),$ (7)
-$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\tilde{\mathbf{s}}-\mathbf{r}^{(k)}\right\|_{2}^{2}+\lambda\|\boldsymbol{\Psi}\tilde{\mathbf{s}}\|_{1}.$ (8)
-where $\boldsymbol{\Psi}$ denotes the transform matrix, $k$ the iteration index, and $\rho$ the step size. The second step represents a proximal mapping:
-$\operatorname{prox}_{\lambda\phi}(\mathbf{r})=\arg\min_{\tilde{\mathbf{s}}}\frac{1}{2}\|\tilde{\mathbf{s}}-\mathbf{r}\|_{2}^{2}+\lambda\phi(\tilde{\mathbf{s}}).$ (9)
+$\tilde{\mathbf{s}}^{(k)}$ $=\underset{\tilde{\mathbf{s}}}{\arg\min}\frac{1}{2}\left\|\tilde{\mathbf{s}}-\mathbf{r}^{(k)}\right\|_{2}^{2}+\lambda\|\bm{\Psi}\tilde{\mathbf{s}}\|_{1},$ (8)
+where $\bm{\Psi}$ denotes the transform matrix, $k$ the iteration index, and $\rho$ the step size. The second step represents a proximal mapping:
+[…]
 While ISTA with orthogonal transforms (e.g., wavelets) has efficient solutions, it faces challenges with complex transforms and requires numerous iterations.
@@
 The overall framework consists of multiple cascaded stages.
-Each stage contains three main components: a dual-branch dynamic transform module ($\mathcal{F}_{d}^{(k)}$) for feature extraction, a dynamic threshold module ($\Theta_{d}^{(k)}$) for feature refinement, and an inverse transform module ($\tilde{\mathcal{F}_{d}}^{(k)}$) for reconstruction.
+Each stage contains three main components: a dual-branch dynamic transform module ($\mathcal{F}_{d}^{(k)}$) for feature extraction, a dynamic threshold module ($\Theta_{d}^{(k)}$) for feature refinement, and an inverse transform module ($\tilde{\mathcal{F}}^{(k)}$) for reconstruction.
 4.2 DISTA-Net: A Dynamic Framework
-To integrate ISTA with network-based sparse representation, DISTA-Net maps each ISTA iteration to a network layer.
-By stacking these layers and enabling end-to-end training, the network optimizes parameters automatically, improving upon traditional ISTA’s manually designed parameters while maintaining interpretability.
-Building upon ISTA-Net’s framework for compressed sensing, we introduce two key improvements.
+Algorithm 1 DISTA-Net
+0: CSIST image $\mathbf{z}$, steering matrix $\mathbf{G}(\mathbf{x},\mathbf{y})$, initial matrix $Q_{\text{init}}$, number of stages NN, step size $\{\rho^{(k)}\}_{k=1}^{N}$
+[…]
+[…]
+2:for $k=1$ to NN do
+3:  $\mathbf{r}^{(k)}\leftarrow\tilde{\mathbf{s}}^{(k-1)}-\rho^{(k)}\mathbf{G}^{\top}(\mathbf{G}\tilde{\mathbf{s}}^{(k-1)}-\mathbf{z})$
+4:  $\acute{\mathbf{s}}^{(k)}\leftarrow\mathcal{F}_{d}(\tilde{\mathbf{s}}^{(k-1)},\mathbf{r}^{(k)})$
+5:  $\theta_{d}^{(k)}\leftarrow\mathrm{DTG}^{(k)}(\acute{\mathbf{s}}^{(k)})$
+6:  $\tilde{\mathbf{s}}^{(k)}\leftarrow\tilde{\mathcal{F}}(\operatorname{Soft}(\mathcal{F}_{d}(\mathbf{r}^{(k)}),\theta_{d}^{(k)}))$
+7:end for
+The details of our DISTA-Net are illustrated in Fig.
+4.
+Building upon ISTA-Net’s framework, we introduce two key improvements.
 First, we design a data-adaptive nonlinear transformation function $\mathcal{F}_{d}(\cdot)$ that maps images to richer dimensional representations while emphasizing significant image regions.
@@
 where $\operatorname{Soft}(\cdot,\theta_{d})$ denotes soft-thresholding with learnable parameter $\theta_{d}$, and $k$ is the stage index.
-The left inverse $\tilde{\mathcal{F}}$ satisfies ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with a structure symmetric to $\mathcal{F}_{d}(\cdot)$.
-This yields:
+The left inverse $\tilde{\mathcal{F}}$ satisfies $\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, without requiring structural symmetry to $\mathcal{F}_{d}(\cdot)$, yielding:
 $\tilde{\mathbf{s}}^{(k)}=\tilde{\mathcal{F}}(\operatorname{Soft}(\mathcal{F}_{d}(\mathbf{r}^{(k)}),\theta_{d})).$ (12)
@@
 4, DISTA-Net comprises $N$ stages.
-Each stage $k$ ($k>1$) contains three components: $\mathcal{F}_{d}^{(k)}$, $\operatorname{Soft}(\cdot,\theta_{d}^{(k)})$, and ~$\tilde{\mathcal{F}}^{(k)}$.
+Each stage $k$ ($k>1$) contains three components: $\mathcal{F}_{d}^{(k)}$, $\operatorname{Soft}(\cdot,\theta_{d}^{(k)})$, and $\tilde{\mathcal{F}}^{(k)}$.
 The input $\mathbf{r}^{(k)}$ is derived from $\tilde{\mathbf{s}}^{(k-1)}$ via Eq.
 (7), processed through these components sequentially to generate $\tilde{\mathbf{s}}^{(k)}$ for the next iteration.
+Overall, the proposed DISTA-Net can be referenced in Algorithm  1.
 Dynamic Transform.
 While trainable non-linear transformations overcome limitations of handcrafted methods, their fixed post-training parameters result in static transformation patterns.
-We address this by introducing a dual-branch Dynamic Transform module $\mathcal{F}_{d}^{(k)}$ at the $k$-th stage (Fig.
-4).
+We address this by introducing a dual-branch Dynamic Transform module $\mathcal{F}_{d}^{(k)}$ at the $k$-th stage.
 To handle the sensitivity of sparse image $\tilde{\mathbf{s}}^{(k-1)}$ perturbations, we design an auxiliary branch that guides $\mathbf{r}^{(k)}$ through a dynamic convolutional kernel.
@@
 The module first processes $\tilde{\mathbf{s}}^{(k-1)}$ through a fully connected network to generate a weight vector:
-$W=f(\tilde{\mathbf{s}}^{(k-1)})$ (13)
-The Dynamic Conv module(Fig. 4) applies this weight vector as an adaptive convolutional kernel to $\mathbf{r}^{(k)}$:
-$w_{r}=C(W,\mathbf{r}^{(k)})$ (14)
+$W=f(\tilde{\mathbf{s}}^{(k-1)}).$ (13)
+The Dynamic Conv module applies this weight vector as an adaptive convolutional kernel to $\mathbf{r}^{(k)}$:
+$w_{r}=C(W,\mathbf{r}^{(k)}).$ (14)
 The final output combines a Conv-ReLU-Conv branch with the sigmoid-activated auxiliary branch:
@@
 Dynamic Soft-Thresholding.
-Unlike ISTA-Net’s fixed threshold $\theta$, we propose a Dynamic Soft-Thresholding module that adapts $\theta_{d}$ based on input image information.
+Unlike ISTA-Net’s fixed threshold $\theta$, we propose a Dynamic Thresholding Generator(DTG) module that adapts $\theta_{d}$ based on input image information.
 This approach better handles densely overlapped targets and spatial context variations.
 As shown in Fig.
-4, the module employs dual convolutional layers to capture multi-scale features as in Li et al.
-(2024a, 2019a).
+4, the module employs dual convolutional layers to capture multi-scale features as in [16, 12].
 The $\mathcal{F}_{d}^{(k)}$ output passes through two $3\times 3$ convolutions, generating feature maps $\tilde{U}_{1}$ and $\tilde{U}_{2}$.
@@
 Spatial relationships are captured through parallel pooling operations:
-$SA_{\text{avg}}=P_{\text{avg}}(\tilde{U}),\quad SA_{\text{max}}=P_{\text{max}}(\tilde{U}).$ (16)
+[…]
 The pooled features are processed through the convolution:
@@
 $(\tilde{SA})_{i}=\text{sigmoid}((\hat{SA})_{i}).$ (18)
-The dynamic threshold $\theta_{d}$ is then computed by combining these masks with multi-scale feature maps through a final convolution.
+The dynamic threshold $\theta_{d}$ is then computed by combining these masks with multi-scale feature maps through a final convolution $C(\cdot)$:
+$\theta_{d}=C(\sum_{i=1}^{N}{(\tilde{SA})_{i}\cdot\tilde{U}_{i}}).$ (19)
 Initialization and Learning Objectives.
@@
 $\mathcal{L}_{\text{constraint}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\sum_{k=1}^{N}\|\tilde{\mathcal{F}}^{(k)}(\mathcal{F}_{d}^{(k)}(s_{i}))-s_{i}\|_{2}^{2}.$ (22)
-Here, $\mathcal{L}_{\text{discrepancy}}$ computes the MSE between reconstructed $\tilde{s}_{i}^{(N)}$ and ground truth $s_{i}$, while $\mathcal{L}_{\text{constraint}}$ enforces multi-stage identity constraints through ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with $\gamma=0.01$ balancing these objectives.
+Here, $\mathcal{L}_{\text{discrepancy}}$ computes the MSE between reconstructed $\tilde{s}_{i}^{(N)}$ and ground truth $s_{i}$, while $\mathcal{L}_{\text{constraint}}$ enforces multi-stage identity constraints through $\tilde{\mathcal{F}}^{(k)}\circ\mathcal{F}_{d}^{(k)}\approx\mathbf{I}$, with $\gamma=0.01$ balancing these objectives.
 5 Experiments
@@
                                           Traditional Optimization
-ISTA Daubechies et al. (2004) - - 7.46 0.01 0.31 2.39 9.46 25.14 - -
+ISTA [7] - - 7.46 0.01 0.31 2.39 9.46 25.14 - -
 Image Super-Resolution
-ACTNet Zhang et al. (2023) 46.212M 62.80G 45.61 0.38 7.46 41.13 83.12 95.95 35.4526 99.70
-CTNet Lim et al. (2017) 0.400M 2.756G 45.11 0.38 7.53 40.39 82.11 95.14 35.1499 99.70
-DCTLSA Zeng et al. (2023) 0.865M 13.69G 44.51 0.39 7.35 39.35 81.15 94.34 34.6314 99.65
-EDSR Lim et al. (2017) 1.552M 12.04G 45.32 0.33 7.07 40.58 83.24 95.41 35.3724 99.71
-EGASR Qiu et al. (2023) 2.897M 17.73G 45.51 0.42 8.03 41.32 85.71 95.08 34.5681 99.66
-FeNet Wang et al. (2022) 0.348M 2.578G 45.77 0.42 8.19 42.13 83.30 94.80 34.1531 99.66
-RCAN Zhang et al. (2018a) 1.079M 8.243G 45.87 0.42 7.96 41.81 83.61 95.57 35.2119 99.69
-RDN Zhang et al. (2018b) 22.306M 173.0G 45.81 0.35 7.11 41.07 84.07 96.43 36.4686 99.74
-SAN Dai et al. (2019) 4.442M 34.05G 45.95 0.36 7.35 41.17 84.32 96.57 36.5037 99.74
-SRCNN Dong et al. (2015) 0.019M 1.345G 29.06 0.23 4.10 21.65 49.95 69.39 28.7608 98.44
-SRFBN Li et al. (2019b) 0.373M 3.217G 46.05 0.43 8.31 42.83 83.72 94.95 34.0174 99.68
-HAN Niu et al. (2020) 64.342M 495.0G 45.70 0.39 7.46 40.90 83.61 96.17 35.2703 99.71
-HiT-SIR Zhang et al. (2024) 0.769M 7.357G 44.51 0.40 7.16 39.55 80.94 94.50 34.7400 99.70
-HiT-SNG Zhang et al. (2024) 0.952M 13.324G 45.01 0.39 7.34 40.19 81.98 95.17 35.1390 99.71
-HiT-SRF Zhang et al. (2024) 0.844M 7.945G 44.71 0.40 7.71 40.03 80.94 94.49 34.5449 99.67
+ACTNet [45] 46.212M 62.80G 45.61 0.38 7.46 41.13 83.12 95.95 36.8879 0.9854
+CTNet [30] 0.400M 2.756G 45.11 0.38 7.53 40.39 82.11 95.14 35.6154 0.9833
+DCTLSA [38] 0.865M 13.69G 44.51 0.39 7.35 39.35 81.15 94.34 35.2259 0.9832
+EDSR [19] 1.552M 12.04G 45.32 0.33 7.07 40.58 83.24 95.41 36.4349 0.9816
+EGASR [25] 2.897M 17.73G 45.51 0.42 8.03 41.32 85.71 95.08 35.4549 0.9745
+FeNet [31] 0.348M 2.578G 45.77 0.42 8.19 42.13 83.30 94.80 36.4507 0.9864
+RCAN [43] 1.079M 8.243G 45.87 0.42 7.96 41.81 83.61 95.57 36.5860 0.9831
+RDN [44] 22.306M 173.0G 45.81 0.35 7.11 41.07 84.07 96.43 36.1883 0.9840
+SAN [2] 4.442M 34.05G 45.95 0.36 7.35 41.17 84.32 96.57 37.1805 0.9848
+SRCNN [8] 0.019M 1.345G 29.06 0.23 4.10 21.65 49.95 69.39 29.4140 0.9566
+SRFBN [17] 0.373M 3.217G 46.05 0.43 8.31 42.83 83.72 94.95 34.2945 0.9815
+HAN [23] 64.342M 495.0G 45.70 0.39 7.46 40.90 83.61 96.17 36.6048 0.9826
+HiT-SNG [42] 0.952M 13.324G 45.01 0.39 7.34 40.19 81.98 95.17 35.4038 0.9805
 Deep Unfolding
-ISTA-Net Zhang and Ghanem (2018) 0.171M 12.77G 45.16 0.41 7.71 40.57 82.58 94.53 33.9215 99.68
-ISTA-Net+ Zhang and Ghanem (2018) 0.337M 24.33G 46.06 0.42 7.66 41.58 84.46 96.17 36.0892 99.72
-LAMP Daubechies et al. (2004) 2.126M 0.278G 14.22 0.05 1.11 7.31 21.56 41.06 27.8299 96.89
-LIHT Daubechies et al. (2004) 21.10M 1.358G 10.35 0.06 0.92 4.99 14.74 30.5 27.5107 96.42
-LISTA Daubechies et al. (2004) 21.10M 1.358G 30.13 0.25 4.13 22.29 51.18 72.82 29.8936 99.12
-FISTA-Net Xiang et al. (2021) 0.074M 18.96G 44.66 0.45 7.68 39.74 81.24 94.19 35.7519 99.67
-TiLISTA Daubechies et al. (2004) 2.126M 0.278G 14.95 0.06 1.23 7.72 22.50 46.23 27.7038 97.40
-$\star$ DISTA-Net (Ours) 2.179M 35.10G 46.74 0.38 7.54 42.44 86.18 97.14 37.8747 99.79
+ISTA-Net [40] 0.171M 12.77G 45.16 0.41 7.71 40.57 82.58 94.53 35.6733 0.9862
+ISTA-Net+ [40] 0.337M 24.33G 46.06 0.42 7.66 41.58 84.46 96.17 38.4998 0.9887
+LAMP [7] 2.126M 0.278G 14.22 0.05 1.11 7.31 21.56 41.06 27.8181 0.9405
+LIHT [7] 21.10M 1.358G 10.35 0.06 0.92 4.99 14.74 30.50 27.5657 0.9198
+LISTA [7] 21.10M 1.358G 30.13 0.25 4.13 22.29 51.18 72.82 30.1285 0.9537
+FISTA-Net [34] 0.074M 18.96G 44.66 0.45 7.68 39.74 81.24 94.19 38.4773 0.9872
+TiLISTA [7] 2.126M 0.278G 14.95 0.06 1.23 7.72 22.50 46.23 27.7790 0.8895
+$\star$ DISTA-Net (Ours) 2.179M 35.10G 46.74 0.38 7.54 42.44 86.18 97.14 38.3825 0.9887
 Table 1: Comparison with SOTA methods on the CSIST-100K dataset.
@@
 Training.
-Using images from the CSIST dataset as input, we employ the method mentioned in  4.1 to perform sub-pixel division with a sampling grid ratio of $c$, generating an unmixed high-resolution grid as ground truth.
-Specifically, for the $i$-th target with parameters $(x_{i},y_{i},g_{i})$, the intensity $g_{i}$ is assigned to the pixel at position $\left(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2}\right)$, with other pixels set to zero.
-With an appropriate selection of c, each target is resolved as a distinct point in the unmixed grid, enabling precise spatial separation.
+Using CSIST-100K images as input, we apply Sec.
+4.1’s method to perform sub-pixel division with a sampling grid ratio of $c$.
+This generates an unmixed high-resolution grid as ground truth, where for each target $(x_{i},y_{i},g_{i})$, the intensity $g_{i}$ is assigned to the pixel at position $\left(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2}\right)$ while other pixels remain zero.
+The selected $c$ value ensures each target appears as a distinct point, enabling accurate spatial separation.
 Testing.
-We project the unmixed high-resolution grid back to the original 11×11 coordinate space to calculate the CSO-mAP against the ground truth target positions.
-Post-processing is applied to the model’s predicted images, where pixels with intensity values above a threshold of 50 are retained as predicted targets.
-For the i-th predicted target, its coordinates are computed as $\left(\frac{x_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c},\frac{y_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c}\right)$.
-To provide a more comprehensive assessment of the model’s performance, we also compute PSNR and SSIM metrics by directly comparing the predicted high-resolution images with the sub-pixel divided ground truth images.
-Hyperparameters.
-We set $c=3$ as the baseline.
-The network is trained on randomly selected batches of 64 images, with the DIST-Net configured with $N=6$ stages and the Dynamic Transform branch coefficient $(1-\alpha)$ set to 0.3.
-5.2 Comparison with SOTA Methods
+We (1) apply post-processing with an intensity threshold of 50 to identify predicted targets; (2) project the unmixed grid back to the original $11\times 11$ space via $\left(\frac{x_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c},\frac{y_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c}\right)$.
+These mapped coordinates are then compared with ground truth target positions to compute CSO-mAP.
+Additionally, we calculate PSNR and SSIM by directly comparing the full predicted and ground truth high-resolution images.
+Hyperparameters. Our configuration uses: $c=3$ (baseline grid ratio), batch size of 64, DISTA-Net with 6 stages ($N=6$), and Dynamic Transform branch coefficient $(1-\alpha)=0.3$.
+5.2 Comparison with State-of-the-Art Methods
 Experimental Results.
 Table  1 compares DISTA-Net with traditional optimization (ISTA), image super-resolution, and deep unfolding methods on the CSIST-100K dataset, evaluated by computational efficiency (#P/FLOPs), localization accuracy (CSO-mAP), and image quality (PSNR/SSIM).
-For localization accuracy, we adopt CSO-mAP with different distance thresholds (from AP-05 to AP-25), where a smaller threshold indicates stricter localization precision requirement.
+For localization accuracy, we adopt CSO-mAP with different distance thresholds (from AP-05 to AP-25), where a smaller threshold indicates a stricter localization precision requirement.
 For instance, AP-05 evaluates the detection accuracy within 0.05-pixel distance, representing an extremely high precision demand.
-Under practical scenarios where 0.25-pixel precision (AP-25) is sufficient, DISTA-Net achieves 97.14% accuracy, outperforming most existing methods.
-Even under more stringent conditions (AP-15 and AP-10), our method maintains competitive performance with 42.44% and 7.54% accuracy respectively, demonstrating its robust localization capability across different precision requirements.
+Considering the inherent error of 0.236 pixel width at subpixel division factor $c=3$, AP-20 and AP-25 (with 0.20 and 0.25 precision requirements respectively) serve as primary performance benchmarks.
+DISTA-Net achieves remarkable accuracy rates of 86.18% and 97.14% accuracy, outperforming most existing methods.
+On the mAP metric reflecting average CSIST unmixing performance, our method maintains a leading advantage with 46.74% accuracy, demonstrating its robust localization capability across different precision requirements.
 In terms of model efficiency, DISTA-Net achieves these results with moderate computational costs (2.179M parameters and 35.103G FLOPs), showing better efficiency compared to methods like ACTNet (46.212M, 62.798G) and HAN (64.342M, 0.495T).
-Additionally, DISTA-Net demonstrates superior image quality with the highest PSNR (37.8747) and SSIM (99.79) scores among all compared methods, indicating its excellent capability in preserving image details while performing target localization.
+Additionally, DISTA-Net achieves a leading SSIM score of 0.9887 and a satisfactory PSNR of 38.3825, effectively preserving image details.
 These comprehensive results validate that DISTA-Net achieves an effective balance between computational efficiency, localization accuracy, and reconstruction quality.
-Figure 5: Visual comparison of 3× sub-pixel division reconstruction for scenes containing different numbers of adjacent targets.
-The red boxes highlight regions where targets exhibit significant sub-pixel characteristics.
 Visual comparison.
 Fig.
-5 compares reconstruction results with 3× sub-pixel division across methods for scenes containing 3–5 adjacent targets, assessing performance in dense multi-target scenarios.
-Existing methods struggle to reconstruct closely-spaced targets, exhibiting blurred boundaries or merged detections, with degradation worsening for higher target counts (e.g., five-target cases).
+5 compares reconstruction results with $3\times$ sub-pixel division across methods for scenes containing $3\sim 5$ targets, assessing performance in dense multi-target scenarios.
+Existing methods struggle to reconstruct closely spaced targets, exhibiting blurred boundaries or merged detections, with degradation worsening for higher target counts (e.g., five-target cases).
 In contrast, DISTA-Net preserves both target quantity and sub-pixel positions while maintaining sharp boundaries and accurate spatial distributions, even under extreme density.
+Figure 5: Visual comparison of $3\times$ sub-pixel division reconstruction for scenes containing different numbers of closely-spaced infrared small targets.
+The red boxes highlight regions where targets exhibit significant sub-pixel characteristics.
 5.3 Ablation Study
-Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP
-  ISTA-Net 0.171M 12.766G 45.16
-DISTA-Net w/o DT 0.171M 39.544G 46.32
-DISTA-Net w/o Thres. 5.111M 92.62G 46.17
-$\star$ DISTA-Net (Ours) 2.179M 35.103G 46.74
+Method CSO-mAP AP-05 AP-10 AP-15 AP-20 AP-25
+  ISTA-Net [40] 45.16 0.41 7.71 40.57 82.58 94.53
+DISTA-Net w/o DT 46.32 0.34 6.83 40.76 86.18 97.50
+DISTA-Net w/o Thres. 46.17 0.44 7.77 42.18 84.67 95.79
+$\star$ DISTA-Net (Ours) 46.74 0.38 7.54 42.44 86.18 97.14
 Table 2: The effect of different components.
@@
 Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP
+mAP AP-10 AP-15
                                                     c=5
-ISTA-Net Zhang and Ghanem (2018) 0.171M 39.544G 66.90
-ISTA-Net+ Zhang and Ghanem (2018) 0.225M 48.158G 68.50
-CFGN Dai et al. (2023a) 0.538M 4.122G 67.95
-$\star$ DISTA-Net (Ours) 5.153M 102.4G 69.58
+ISTA-Net  [40] 0.171M 39.544G 66.90 56.73 87.26
+ISTA-Net+  [40] 0.225M 48.158G 68.50 57.96 89.52
+CFGN  [3] 0.538M 4.122G 67.95 58.08 88.35
+$\star$ DISTA-Net (Ours) 5.153M 102.4G 69.58 60.95 90.85
 c=7
-ISTA-Net Zhang and Ghanem (2018) 0.171M 89.51G 71.19
-ISTA-Net+ Zhang and Ghanem (2018) 0.225M 103.0G 71.09
-CFGN Dai et al. (2023a) 0.548M 4.202G 70.38
-$\star$ DISTA-Net (Ours) 6.409M 142.3G 72.84
+ISTA-Net  [40] 0.171M 89.51G 71.19 76.45 84.16
+ISTA-Net+  [40] 0.225M 103.0G 71.09 74.90 84.90
+CFGN  [3] 0.548M 4.202G 70.38 73.88 83.97
+$\star$ DISTA-Net (Ours) 6.409M 142.3G 72.84 78.47 86.09
 Table 3: Comparison of methods across different Sampling Grids on the CSIST-100K dataset.
 Model Performance and Sampling Grid.
-To investigate the impact of sampling grid ratio on model performance and computational efficiency, we conduct experiments with other different sampling grid ratios (c = 5, and 7) across various methods.
-Table 3 presents a comprehensive comparison of model performance under different sampling configurations.
-As the sampling grid ratio increases, we observe a general trend of improved detection performance across all methods, which is attributed to the higher precision in target position representation.
-DISTA-Net consistently maintains a leading advantage in positioning accuracy, outperforming other methods at all configurations.
-However, this performance gain comes at the cost of substantially increased computational complexity.
-Notably, DISTA-Net demonstrates better efficiency in utilizing the increased resolution.
-Compared to ISTA-Net and ISTA-Net+, which show similar trends but lower performance, our method achieves higher accuracy gains with the same increase in sampling ratio.
-Based on these observations, we suggest selecting an appropriate sampling grid ratio according to real-world requirements to achieve a good balance between detection accuracy and computational efficiency.
-Hyperparameters Analysis. We analyzed model performance versus stage numbers and dynamic branch coefficients in Supplementary, demonstrating robust architectural and parametric design.
+We evaluate sampling ratios $c=5$ and $7$ across methods, analyzing their impact on performance and efficiency.
+As the sampling grid ratio increases, all methods show improved detection performance due to better target positioning precision.
+DISTA-Net consistently outperforms other methods across all configurations.
+While ISTA-Net and ISTA-Net+ exhibit similar trends with lower overall performance, our proposed method achieves greater accuracy improvements for equivalent increases in sampling ratio.
+This advantage is particularly pronounced in the AP-10 ($c=5$) and AP-15 ($c=7$) metrics (see Table 3 for complete results).
+However, these performance gains are accompanied by substantially increased computational complexity.
+We recommend selecting an appropriate sampling grid ratio that aligns with specific application requirements to achieve an optimal balance between detection accuracy and computational efficiency.
+Ours vs Super-Resolution + Detector Pipeline.
+The unmixing stage typically serves as a refinement step following the detection phase.
+For experimental rigor, we conducted experiments implementing the “SR + Detector” pipeline.
+We used YOLOv11 as the detector following leading SR methods (retrained on our IR data with unmixing GT for point-source super-resolution).
+The results demonstrate DISTA-Net’s continued advantage: DISTA-Net + YOLOv11 achieved a CSO-mAP of 47.82, outperforming SRFBN + YOLOv11 (45.74) and CFGN + YOLOv11 (46.71).
+Visual analysis reveals that both conventional SR methods and our unmixing approach generate high-resolution images with well-resolved peaks, enabling effective target separation through simple thresholding.
+This demonstrates the dominant role of the unmixing stage in this task.
+Hyperparameters Analysis. We analyzed model performance versus stage numbers and dynamic branch coefficients in Supplementary, demonstrating robust design.
 6 Conclusion
-In this work, we address the challenging problem of super-resolving closely-spaced infrared small targets, where accurate target counting, sub-pixel positioning, and radiation intensity measurement pose significant challenges.
-We present DISTA-Net, a deep learning framework that incorporates dynamic spatial attention mechanisms to reformulate traditional sparse reconstruction.
-Our comprehensive experimental results demonstrate that DISTA-Net achieves superior performance in both sub-pixel target detection accuracy and image reconstruction quality.
+In this paper, we present the Dynamic Iterative Shrinkage Thresholding Network (DISTA-Net) to address CSIST unmixing task, which features adaptive generation of both convolution weights and thresholding parameters.
+Extensive experiments demonstrate that DISTA-Net achieves superior performance in both sub-pixel target detection accuracy and image reconstruction quality.
 To advance research in this domain, we introduce the CSIST dataset, CSO-mAP metric, and GrokCSO toolkit.
-Supplementary
-1 Multi-target Imaging
-Figure 6: Multi target imaging, the imaging of distant targets on the image plane can be regarded as the diffusion of energy by point source targets through PSF function.
-Multi target imaging is the cumulative response of superimposed point sources.
-When targets are at considerable distances, they can be approximated as point sources, with their energy distribution on the sensor’s focal plane characterized by Point Spread Functions (PSFs) that approximate two-dimensional Gaussian functions.
-When multiple targets are present in close proximity, the detector receives a combined energy response that is the linear sum of individual PSFs .
-This superposition principle is fundamental to the super-resolution unmixing process, where the stacked response must be decomposed to identify and characterize closely-spaced infrared targets, as illustrated in Fig.
-6
-2 CSO-mAP Metric
-Once the predictions are classified as TP or FP, a binary list is constructed following the common practice in COCO, where a TP prediction is indicated by 1 and an FP prediction by 0.
-This binary list serves as the foundation for generating the Precision-Recall (PR) curve.
-By varying the confidence threshold for positive predictions, a series of Precision and Recall values are obtained, forming the PR curve.
-The Average Precision (AP), calculated as the area under the PR curve, provides a comprehensive assessment of the model’s performance across various confidence thresholds, effectively summarizing the trade-off between Precision and Recall.
-Finally, we introduce the CSO mean Average Precision (CSO-mAP), which averages the AP across different distance thresholds $\delta_{k}$, offering a standardized metric for comparing the performance of models in the CSIST unmixing task.
-3 GrokCSO Toolkit
-Recognizing the void in specialized, accessible tools for this domain, we introduce GrokCSO, a comprehensive, open-source toolkit architected to enhance the reconstruction of closely spaced infrared small targets.
-While generic computer vision fields benefit from a plethora of object detection toolkits like MMDetection and GluonCV, the specificity of infrared small target unmixing has long necessitated a dedicated platform.
-The absence of such a platform has fragmented research efforts, impeding the reproducibility of experiments and the comparative analysis of algorithms.
-Developed atop the robust PyTorch framework, GrokCSO is meticulously designed to address the unique challenges inherent to closely spaced infrared small target unmixing.
-It stands distinguished in its offerings:
-  1. 1.
-    Pre-trained Models and Reproducibility: GrokCSO equips researchers with an arsenal of pre-trained models, complete with training scripts and logs of leading-edge algorithms.
-This repository of resources not only fosters reproducibility but also facilitates nuanced comparisons across diverse algorithmic approaches.
-  2. 2.
-    Tailored Flexibility and Evaluative Rigor: With an extensive suite of adaptable backbones and necks, GrokCSO accommodates a broader spectrum of computational strategies.
-The toolkit incorporates specialized dataset loaders, cutting-edge attention mechanisms, and versatile data augmentation workflows.
-Moreover, GrokCSO introduces evaluation metrics specifically calibrated for the intricacies of closely spaced infrared small target unmixing.
-These metrics honor the distinct nature of CSO challenges.
-4 Hyperparameters Analysis
-Figure 7: CSO-mAP performance under different Stage Numbers and Branch coefficients.
-We also conduct experiments to analyze the relationship between model performance, the number of stages, and the coefficient of the dynamic transform branch in Supplementary, validating its architectural and parametric robustness.
-First, to determine an appropriate number of stages, we trained the network with stage numbers $K=2,4,6,8,$ and $10$.
-The performance curves are recorded in Figure 7.
-As the number of stages increases from 2 to 6, the CSO-mAP score improves, reaching its peak (46.74%) at 6 stages.
-Beyond this point, additional stages result in diminishing returns, likely due to increased model complexity.
-Therefore, selecting 6 stages provides an optimal balance between accuracy and efficiency.
-Second, we adjusted the proportion of the dynamic transform branch relative to the total contribution of both branches and observed the performance changes.
-Starting from 0% coefficient, performance improves as the influence of the dynamic branch increases, peaking at a 30% coefficient.
-The model maintains stable performance across coefficient settings from 50% to 90%, demonstrating robustness.
-However, at 100% coefficient, performance drops sharply to 30.62 (not labeled in the figure), confirming that integrating both branches is crucial for optimal feature representation.
-5 Initialization and Learning Objectives.
-Initialization.
-Given a dataset $\{(z_{i},s_{i})\}_{i=1}^{M}$, where $z_{i}$ is an image of overlapping targets and $s_{i}$ is the corresponding super-resolved image, we calculate high-resolution target coordinates in $s_{i}$ using the target information $(x_{i},y_{i},g_{i})$ and a scaling factor $c$.
-For instance, the grayscale value $g_{i}$ at position $(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2})$ in the high-resolution image is used to generate $s_{i}$.
-Let $Z=[z_{1},\ldots,z_{M}]$ and $S=[s_{1},\ldots,s_{M}]$. The matrix $Q_{\text{init}}$ for initializing $\tilde{s}^{(0)}$ is computed as:
-$Q_{\text{init}}=\arg\min_{Q}\|QZ-S\|_{F}^{2}=SZ^{T}(ZZ^{T})^{-1}.$
-Thus, $\tilde{s}^{(0)}$ is given by:
-$\tilde{s}^{(0)}=Q_{\text{init}}Z.$
-Loss Function.
-To ensure that the unmixed image $\tilde{s}$ is as close as possible to the ground truth image $s$ while maintaining the structure $\tilde{F}(\cdot)\circ F(\cdot)=\mathbf{I}$, we designed the following end-to-end training loss function for DISTA-Net, with a training dataset of size $M$, $N$ stages, and image size $N_{s}$:
-$\mathcal{L}=\mathcal{L}_{\text{discrepancy}}+\gamma\mathcal{L}_{\text{constraint}},$
-where:
-$\mathcal{L}_{\text{discrepancy}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\|\tilde{s}_{i}^{(N)}-s_{i}\|_{2}^{2},$
-$\mathcal{L}_{\text{constraint}}=\frac{1}{MN_{s}}\sum_{i=1}^{M}\sum_{k=1}^{N}\|\tilde{F}^{(k)}(F^{(k)}(s_{i}))-s_{i}\|_{2}^{2}.$
-Here, $L_{\text{discrepancy}}$ measures the mean squared error (MSE) between the super-resolved image $\tilde{s}_{i}^{(N)}$ and the ground truth image $s_{i}$.
-$L_{\text{constraint}}$ enforces the structural constraint by ensuring that the composition of $\tilde{F}^{(k)}$ and $F^{(k)}$ approximates the identity transformation for each stage $k$.
-$\gamma$ is a parameter that balances the discrepancy and constraint terms.
-This loss function aims to balance the accuracy of the unmixed images with the structural integrity of the transformation functions.
+Acknowledgement This research is supported by the NSFC (NO.62206133, 62301261, 62206134, U24A20330, 62361166670, 62225604) and the Shenzhen Science and Technology Program (JCYJ20240813114237048).
+Computation is supported by the Supercomputing Center of Nankai University.